OpenStudy (anonymous):

Emma already has read 6 of 20 books on her summer reading list. What percent of the books on her list has she read already? (2 points

OpenStudy (anonymous):

30%

hero (hero):

part/whole = %/100 6/20 = x/100 20x = 600 x = 600/20 x = 30